OpenStudy (anonymous):

If sin θ = 3 over 7 and tan θ < 0, what is the value of cos θ?

OpenStudy (phi):

I would first sketch a triangle  sin is opposite over hypotenuse so 3/7 means opposite side is 3 and hypotenuse is 7 what is the "x" side ? we would use pythagoras to find x x^2 + y^2 = h^2 x^2 + 3^2 = 7^2 can you find x ?

OpenStudy (anonymous):

radical 10?

OpenStudy (phi):

first simplify the equation to x^2 + 9= 49 now add -9 to both sides

OpenStudy (anonymous):

58??

OpenStudy (phi):

49-9 = ?

OpenStudy (phi):

you get x^2 = 40 now take the square root of both sides

OpenStudy (anonymous):

would it be 2 radical 10?

OpenStudy (phi):

once we know x, we can find cos however, the angle could be in any of 4 quadrants. but we know the sin is 3/7 is positive. the hypotenuse is always positive. but we would also have these: